# Quantum chaos and the arrow of time

### v10

@@ 3.1 First pass: the arrow of time
$\Delta E_{\beta}=\frac{1}{2}\int_{-\infty}^{\infty}d\omega\;\omega\;|\lambda(\omega)|^{2}\,A_{\beta}(\omega),$ (11)
$A_{\beta}(\omega)=\frac{1}{Z}\left(e^{\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n})),$ (12)
where $Z$ is the partition function. What is remarkable is that the above expression of $\Delta E_{\beta}$ is positive (negative or zero) if the initial temperature is positive (negative or infinite) which is precisely the statement of the second law of thermodynamics in the Kelvin form. The above result is derived using two different techniques in Appendix B. $A_{\beta}(\omega)$ is the spectral function. It is also equal to $-2\,\text{Im}\;\tilde{G}_{R}(\omega)$ where $\tilde{G}_{R}(\omega)$ is the Fourier transform of the retarded Green function $G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle$. So Berry’s conjecture and equivalence of ensemble implies that the quantum chaotic many body systems possess the arrow of time in the thermodynamic limit.
@@
Now let us consider a more general state. Consider a typical state $|\Psi\rangle=\sum_{n}c_{n}|n\rangle$. The $c_{n}$ coefficients are non-zero for a small energy window similar in size to the microcanonical energy window ($dE$ in Appendix C). Now the change in energy after the system is perturbed will have diagonal terms and off-diagonal terms in the $|n\rangle$ basis. We will not be concerned with the various energy scales associated with typical states [19]. Our main motive is to show that the off-diagonal terms average out to zero. The change in energy starting from $|\Psi\rangle$ is
$\Delta E_{\Psi}=\sum_{n}|c_{n}|^{2}\Delta E_{n}+\sum_{n,n^{\prime}\neq n}c^{*}_{n}c_{n^{\prime}}\sum_{m}\Delta E_{nn^{\prime}m},$ (13)
$\sum_{m}\Delta E_{nn^{\prime}m}=\sum_{m}\mathcal{O}_{nm}\mathcal{O}_{mn^{\prime}}\left[\int_{-\infty}^{\infty}dt_{1}dt^{\prime}_{1}\lambda(t_{1})\lambda(t^{\prime}_{1})\,E_{m}e^{it_{1}(E_{n}-E_{m})+it^{\prime}_{1}(E_{m}-E_{n^{\prime}})}\right.\hskip 85.35826pt$
$\left.\hskip 85.35826pt-\int_{-\infty}^{\infty}dt_{1}\int_{-\infty}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\left\{E_{n}e^{it_{2}(E_{n}-E_{m})+it_{1}(E_{m}-E_{n^{\prime}})}+(t_{1}\leftrightarrow t_{2},n\leftrightarrow n^{\prime})\right\}\right],$
where sum over $m$ is over the entire state space. $R_{mn}$ has zero mean so the off-diagonal terms average out to zero. It is worth noting that this will not happen in integrable systems. The (pseudo-) randomness in $R_{mn}$ is at the level of the energy gap between adjacent energy levels which is exponentially small in system size so all the other factors can be effectively taken to be constants compared to the variation of $R_{mn}$. In fact, the energy gap between adjacent energy levels is the smallest energy scale in the system and the Poincare recurrence time is of the order of the inverse of the energy gap. The diagonal terms give the microcanonical average which is the expected result. Averaging of $R_{mn}$ is routinely used in all ETH works, see for example in [19, 21]. We will use it again in the following subsection.
3.2 Second pass: a new constraint on ETH statement
In this subsection, we will do a detailed comparison of (8) and (11) which gives the new constraints - ETH-monotonicity. We use the ETH statement (2) in (11). $|R_{mn}|^{2}$ is replaced by its average value which is equal to one. We substitute $\sum_{m}\to\int_{-\infty}^{\infty}dE_{m}e^{S(E_{m})}$ and $\sum_{n}\to\int_{-\infty}^{\infty}dE_{n}e^{S(E_{n})}$ and convert the integration variables into $\omega^{\prime}=E_{m}-E_{n}$ and $\bar{E}=(E_{m}+E_{n})/2$. (12) becomes
$A_{\beta}(\omega)$ $=$ $(e^{\beta\omega}-1)\int^{\infty}_{-\infty}d\omega^{\prime}d\bar{E}\,e^{S(\bar{E}+\omega^{\prime}/2)+S(\bar{E}-\omega^{\prime}/2)-S(\bar{E})}e^{-\beta\bar{E}-\beta\omega^{\prime}/2}f(\bar{E},\omega)^{2}\delta(\omega-\omega^{\prime})$ (14)
@@ 3.2 Second pass: a new constraint on ETH statement
$=$ $2\sinh(\beta\omega/2)f(\bar{E},\omega)^{2}$
@@
The $\bar{E}$-integral fixes $\bar{E}$ to be the energy corresponding to the inverse temperature $\beta$ while going from the second line to the third line. So, the change in energy starting from the canonical ensemble is
$\Delta E_{\beta}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\sinh(\beta\omega/2)f(\bar{E},\omega)^{2}$ (15)
@@
Now it is clear that $\Delta E_{\beta}$ have the same sign as $\beta$ and is zero when $\beta=0$.
Substituting $\sum_{E_{m}}\to\int_{-\infty}^{\infty}dE_{m}e^{S(E_{m})}$ and changing the integration variable to $\omega=E_{m}-E_{n}$, (8) becomes
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\,e^{S(E_{n}+\omega)-S(E_{n}+\omega/2)}f(E_{n}+\omega/2,\omega)^{2}$ (16)
@@
In a finite system with finite dimensional state space, the entropic factor $e^{S(E_{n}+\omega)-S(E_{n}+\omega/2)}$ will have a peak at $\omega=\omega_{0}$ which has the same sign as $\beta$. We study this entropic factor in detail in Appendix D. We show that the position of the peak of this entropic factor moves away to $\omega\to\pm\infty$ in the thermodynamic limit. For such large systems, the convergence of the above integral is ensured by the $\omega$-dependence of $f(\bar{E},\omega)$ which falls exponentially faster than $e^{-|\beta\omega|/4}$ for large $\omega$. So for large systems, we can perform Taylor expansion of the entropic factor about $\bar{E}=E_{n}$ (similar expansion is used in [21]). This reduces the change in energy to
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\,e^{\beta\omega/2}f(\bar{E}+\omega/2,\omega)^{2}$ (17)
@@
where $\beta$ is the inverse temperature of $\bar{E}=E_{n}$. Note that we have replace $E_{n}$ with $\bar{E}$ in the expression above so that no confusion arises when we compare this expression with (15). The $e^{\beta\omega/2}$ factor is the entropic factor in support of the arrow of time. It comes from the competition of the entropic factor from the sum over the state space and the entropic factor of the off-diagonal term in the ETH statement. So, $f(\bar{E}+\omega/2,\omega)^{2}$ acts like a bare transition amplitude in a quantum chaotic system. Note that the real transition amplitude between energy eigenstates $|n\rangle$ and $|m\rangle$ is simply $|\mathcal{O}_{mn}|^{2}$.
@@
The difference between (17) and (15) is the $\omega/2$ in the first argument of the $f$-function in (15). So, if we perform a Taylor expansion of $f(\bar{E}+\omega/2,\omega)$ about the point $(\bar{E},\omega)$ and consider only the first order term, then we get
@@
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}e^{\beta\omega/2}f(\bar{E},\omega)^{2}\left[1+\frac{\omega}{f(\bar{E},\omega)}\frac{\partial f(\bar{E},\omega)}{\partial\bar{E}}\right]$ (18)
@@
$|\tilde{\lambda}(\omega)|$ and $f(\bar{E},\omega)$ are even functions of $\omega$, so the first term inside the square bracket gives us the thermal value (15). So, the second term inside the square bracket should vanish in the thermodynamic limit. It means that $f(\bar{E},\omega)$ flattens,
[…]
@@
So, the bare transition amplitude is constrained by the arrow of time in a quantum chaotic system. We will present here numerical results for the spin current operator. Numerical results for the kinetic energy operator and the single site spin operator can be found in Appendix E. Figure 3 are plots of the $f$-function as a function of $\bar{E}$ for different system sizes. The flattening of $f(\bar{E})$ is not clear from these plots.
@@
We can identify the scaling of the slope with system size $L$ (for a fixed $L/N$ ratio). For this, we plot the slopes of the linear fits of the $f^{\prime}(\bar{E})/f(\bar{E})$ versus $\bar{E}$ plots for varying $\omega\in[0.1,2.5]$. To calculate the slopes reliably we use the $f^{\prime}(\bar{E})/f(\bar{E})$ versus $\bar{E}$ plots with $\bar{E}\in[-1,1]$. Panel (a) in Figure 5 is plots of the slope as a function of $\omega$. We find that the magnitude of the slope increases slightly with increasing $\omega$. What is more obvious is the flattening of $f$-function with increasing system size. Almost for the entire range of $\omega$, the slope of the largest system size $L=18,N=9$ is the lowest.
Panel (b), (c) and (d) are plots of the slope of the linear fits as a function $\omega$ with different system size scaling. Panel (c) clearly shows that the slopes of the different system sizes $L$ but scaled with $L^{1}$ mostly overlap. This implies that the $f$-function flattens as $L^{1}$. As mathematical expressions,
$\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;\xrightarrow{L\to\infty}\;0,\qquad L\;\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;=\;\text{a constant.}$. (20)
@@
This monotonically increasing behaviour of $f$-function as a function of $|T|$ actually further reinforces the Kelvin statement, meaning the contribution from this variation is such that it gives extra support in making the change in energy $\Delta E$ to have the same sign as the inverse temperature $\beta$.
@@
$\lambda(t)=2\lambda_{0}\Lambda\,\text{sinc}(\Lambda t)$ (21)
$\tilde{\lambda}(\omega)=\lambda_{0}\,\text{rect}\left(\frac{\omega}{2\Lambda}\right)=\lambda_{0}\begin{cases}0,&\quad\text{if}\quad|\omega|>\Lambda\\ \frac{1}{2},&\quad\text{if}\quad|\omega|=\Lambda\\ 1,&\quad\text{if}\quad|\omega|<\Lambda\\ \end{cases}$ (22)
Considering a large enough system size $L$, we can make $\Lambda$ as small as we want without taking it smaller that the random matrix theory (RMT) scale $\sim 1/L^{2}$. We can also ensure that $\Lambda<|T|=1/|\beta|$. Then we can use the expansion of the exponential entropic factor in (17). For small $\Lambda$, we consider only the leading order term which gives us
$\Delta E_{n}=\int_{0}^{\Lambda}d\omega\,|\lambda(\omega)|^{2}\,\omega\left(f(\bar{E}+\omega/2,\omega)^{2}-f(\bar{E}-\omega/2,\omega)^{2}\right)$ (23)
@@
We want to emphasize that we are working in the small $\omega$ region so this argument is not valid for large $\omega$ region. More crucially, there is no reason why the above expression should be either positive or negative, unlike in subsection 3.1 where we invoked Berry’s conjecture. But it is still a highly desirable result otherwise the chaotic system will violate the second law of thermodynamics in this low energy dynamics. Note that this leading $\omega^{1}$ term is absent in the thermodynamic limit because the $f$-function goes to the flat-limit of this monotonic behaviour.
Another motivation for the monotonically increasing behaviour of $f$-function as a function $|T|$ comes from the large $\omega$ behaviour of $f(\beta,\omega)$. As a function of $\omega$, it falls exponentially faster than $e^{-|\beta\omega|/4}$. This is indeed a monotonically increasing function of $|T|=1/|\beta|$ [^5] We thank Anatoly Dymarsky for pointing out this result interpreted in terms of the monotonic behaviour of $f(\beta)$. Again we want to emphasize that this exponential fall-off takes into account only $\omega$-dependence in $f(\beta,\omega)$. There could be other $\beta$-dependence even in the large $\omega$ region. But if we take large enough $\omega$, the other $\beta$-dependence cannot trump the monotonic behaviour. So, at least we have this proof of the monotonic behaviour for large enough $\omega$.
@@
This result along with (20) constitute ETH-monotonicity. We propose that every quantum chaotic many body system will have this property.
Note that temperature is an intensive quantity so the monotonically increasing behaviour will remain intact even in the thermodynamic limit. It can be verified from Figure 7 which are plots of $1/f\partial f/\partial\beta$ as a function of $\beta$ for fixed values of $\omega$. Unlike $f(\bar{E})$, $f(\beta)$ does not flatten as the system size increases. But it is not physically relevant because what appeared in (17) is $f$-function as a function of $\bar{E}$, not as a function of $T$ or $\beta$.

### v12

@@ 3.1 First pass: the arrow of time
$\Delta E_{\beta}=\frac{1}{2}\int_{-\infty}^{\infty}d\omega\;\omega\;|\lambda(\omega)|^{2}\,A_{\beta}(\omega),$ (3.6)
$A_{\beta}(\omega)=\frac{1}{Z}\left(e^{\beta\omega}-1\right)\sum_{n,m}e^{-\beta E_{m}}|\mathcal{O}_{nm}|^{2}\delta(\omega-(E_{m}-E_{n})),\$(3.7)
where $Z$ is the partition function and $\tilde{\lambda}(\omega)=\tilde{\lambda}^{*}(-\omega)$ is the Fourier transform of $\lambda(t)$. What is remarkable is that the above expression of $\Delta E_{\beta}$ is positive (negative or zero) if the initial temperature is positive (negative or infinite) which is precisely the statement of the second law of thermodynamics in the Kelvin form. This is a mathematical identity. The above result is derived using two different techniques in Appendix B. $A_{\beta}(\omega)$ is the spectral function. It is also equal to $2\,\text{Im}\;\tilde{G}_{R}(\omega)$ where $\tilde{G}_{R}(\omega)$ is the Fourier transform of the retarded Green function $G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle$. So Berry’s conjecture and equivalence of ensemble implies that the quantum chaotic many body systems possess the arrow of time in the thermodynamic limit.
@@
Now let us consider a more general state. Consider a typical state $|\Psi\rangle=\sum_{n}c_{n}|n\rangle$. The $c_{n}$ coefficients are non-zero for a small energy window similar in size to the microcanonical energy window ($dE$ in Appendix C). Now the change in energy after the system is perturbed will have diagonal terms and off-diagonal terms in the $|n\rangle$ basis. We will not be concerned with the various energy scales associated with typical states Srednicki_1999 . Our main motive is to show that the off-diagonal terms average out to zero. The change in energy starting from $|\Psi\rangle$ is
$\Delta E_{\Psi}=\sum_{n}|c_{n}|^{2}\Delta E_{n}+\sum_{n,n^{\prime}\neq n}c^{*}_{n}c_{n^{\prime}}\sum_{m}\Delta E_{nn^{\prime}m},$ (3.8)
[…]
$\left.\hskip 85.35826pt-\int_{-\infty}^{\infty}dt_{1}\int_{-\infty}^{t_{1}}dt_{2}\,\lambda(t_{1})\lambda(t_{2})\left\{E_{n}e^{it_{2}(E_{n}-E_{m})+it_{1}(E_{m}-E_{n^{\prime}})}+(t_{1}\leftrightarrow t_{2},n\leftrightarrow n^{\prime})\right\}\right],\$
where sum over $m$ is over the entire state space. $R_{mn}$ has zero mean so the off-diagonal terms average out to zero, so the second sum in (3.8) is zero. It is worth noting that this will not happen in integrable systems. The (pseudo-) randomness in $R_{mn}$ is at the level of the energy gap between adjacent energy levels which is exponentially small in system size so all the other factors can be effectively taken to be constants compared to the variation of $R_{mn}$. In fact, the energy gap between adjacent energy levels is the smallest energy scale in the system. The diagonal terms give the microcanonical average which is the expected result. Averaging of $R_{mn}$ is routinely used in all ETH works, see for example in Srednicki_1999 ; Murthy:2019fgs . We will use it again in the following subsection. We can also consider perturbation with multiple operators satisfying ETH.
$H(t)=H_{0}+\sum_{i}\lambda_{i}(t)\mathcal{O}_{i}$ (3.9)
In this case also, the contribution to the change in energy from the cross-terms $\mathcal{O}_{i}\mathcal{O}_{i^{\prime}}$, with $i\neq i^{\prime}$, will be zero because of $R_{i,mn}$ and $R_{i^{\prime},mn}$.
3.2 Second pass: new constraints on ETH statement
In this subsection, we will do a detailed comparison of (3.3) and (3.6) which gives the new constraints that we are calling ETH-monotonicity. We use the ETH statement (1.2) in (3.6). $|R_{mn}|^{2}$ is replaced by its average value which is equal to one. We substitute $\sum_{m}\to\int_{-\infty}^{\infty}dE_{m}e^{S(E_{m})}$ and $\sum_{n}\to\int_{-\infty}^{\infty}dE_{n}e^{S(E_{n})}$ and convert the integration variables into $\omega^{\prime}=E_{m}-E_{n}$ and $\bar{E}=(E_{m}+E_{n})/2$. (3.7) becomes
$A_{\beta}(\omega)$ $=$ $(e^{\beta\omega}-1)\int^{\infty}_{-\infty}d\omega^{\prime}d\bar{E}\,e^{S(\bar{E}+\omega^{\prime}/2)+S(\bar{E}-\omega^{\prime}/2)-S(\bar{E})}e^{-\beta\bar{E}-\beta\omega^{\prime}/2}f(\bar{E},\omega)^{2}\delta(\omega-\omega^{\prime})$ (3.10)
@@ 3.2 Second pass: new constraints on ETH statement
$=$ $2\sinh(\beta\omega/2)f(\bar{E}_{\beta},\omega)^{2}\$
@@
The $\bar{E}$-integral fixes $\bar{E}$ to be $\bar{E}_{\beta}$, the energy corresponding to the inverse temperature $\beta$, while going from the second line to the third line. In the Taylor expansion of $S(\bar{E}+\omega^{\prime}/2)+S(\bar{E}-\omega^{\prime}/2)-S(\bar{E})$ about $\bar{E}$, the higher order $\omega$-terms are suppressed by powers of the system size. For example, the second order derivative is suppressed by the total heat capacity of the system Murthy:2019fgs . So, the change in energy starting from the canonical ensemble is
$\Delta E_{\beta}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\sinh(\beta\omega/2)f(\bar{E}_{\beta},\omega)^{2}\$(3.11)
@@
Now it is clear that $\Delta E_{\beta}$ have the same sign as $\beta$ and is zero when $\beta=0$. $|\lambda(\omega)|$ and $f(\bar{E},\omega)$ are even functions of $\omega$.
Substituting $\sum_{E_{m}}\to\int_{-\infty}^{\infty}dE_{m}e^{S(E_{m})}$ and changing the integration variable to $\omega=E_{m}-E_{n}$, (3.3) becomes
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\,e^{S(E_{n}+\omega)-S(E_{n}+\omega/2)}f(E_{n}+\omega/2,\omega)^{2}$ (3.12)
@@
In a finite system with finite dimensional state space, the entropic factor $e^{S(E_{n}+\omega)-S(E_{n}+\omega/2)}$ will have a peak at $\omega=\omega_{0}$ which has the same sign as $\beta$. We study this entropic factor in detail in Appendix D. We show that the position of the peak of this entropic factor moves away to $\omega\to\pm\infty$ in the thermodynamic limit. For such large systems, the convergence of the above integral is ensured by the $\omega$-dependence of $f(\bar{E},\omega)$ which falls exponentially faster than $e^{-|\beta\omega|/4}$ for large $\omega$. So for large systems, we can perform Taylor expansion of $S(E_{n}+\omega)-S(E_{n}+\omega/2)$ about $E_{n}$. Again, the higher order derivatives are suppressed by powers of the system size. This reduces the change in energy to
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}\,e^{\beta\omega/2}f(\bar{E}_{\beta}+\omega/2,\omega)^{2}$ (3.13)
@@
where $\beta$ is the inverse temperature of $E_{n}$. Note also that we have replace $E_{n}$ with $\bar{E}_{\beta}$ in the expression above so that no confusion arises when we compare this expression with (3.11). The $e^{\beta\omega/2}$ factor is the entropic factor in support of the arrow of time. It comes from the competition of the entropic factor from the sum over the state space and the entropic factor of the off-diagonal term in the ETH statement. So, $f(\bar{E}_{\beta}+\omega/2,\omega)^{2}$ acts like a bare transition amplitude in a quantum chaotic system. Note that the real transition amplitude between energy eigenstates $|n\rangle$ and $|m\rangle$ is simply $|\mathcal{O}_{mn}|^{2}$. An interesting observation is that $f(\bar{E})$ should be a flat function at $\beta=0$ otherwise $\Delta E_{n}$ will be non-zero which is not expected from the Kelvin statement.
@@
The difference between (3.11) and (3.13) is the $\omega/2$ in the first argument of the $f$-function in (3.11). So, if we perform a Taylor expansion of $f(\bar{E}_{\beta}+\omega/2,\omega)$ about the point $(\bar{E}_{\beta},\omega)$ and consider only the first order term, then we get
@@
$\Delta E_{n}=\int_{-\infty}^{\infty}d\omega\,\omega|\lambda(\omega)|^{2}e^{\beta\omega/2}f(\bar{E}_{\beta},\omega)^{2}\left[1+\left.\left(\frac{\omega}{f(\bar{E},\omega)}\frac{\partial f(\bar{E},\omega)}{\partial\bar{E}}\right)\right|_{\bar{E}=\bar{E}_{\beta}}\right]\$(3.14)
@@
$|\tilde{\lambda}(\omega)|$ and $f(\bar{E},\omega)$ are even functions of $\omega$, so the first term inside the square bracket gives us the thermal value (3.11). Then, the second term inside the square bracket should vanish in the thermodynamic limit. It is a competition with a number, namely one, not a function of $\bar{E}$. It means that $f(\bar{E},\omega)$ flattens,
$\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\quad\xrightarrow[\text{limit}]{\text{Thermodynamic}}\quad 0.$ (3.15)
@@
So, the bare transition amplitude is constrained by the arrow of time in a quantum chaotic system. We will present here numerical results for the spin current operator (an extensive operator). Numerical results for the kinetic energy operator (another extensive operator) and the single site spin operator (an intensive operator) can be found in Appendix E. Figure 3 are plots of the $f$-function as a function of $\bar{E}$ for different system sizes. The flattening of $f(\bar{E})$ is not clear from these plots.
@@
We can identify the scaling of the slope with system size $L$ (for a fixed $L/N$ ratio). For this, we plot the slopes of the linear fits of the $f^{\prime}(\bar{E})/f(\bar{E})$ versus $\bar{E}$ plots for varying $\omega\in[0.1,2.5]$. To calculate the slopes reliably we use the $f^{\prime}(\bar{E})/f(\bar{E})$ versus $\bar{E}$ plots with $\bar{E}\in[-1,1]$. Panel (a) in Figure 5 is plots of the slope as a function of $\omega$. We find that the absolute value of the slope increases slightly with increasing $\omega$. What is more obvious is the flattening of $f$-function with increasing system size. Almost for the entire range of $\omega$, the slope of the largest system size $L=18,N=9$ is the lowest in absolute value. Panel (b), (c) and (d) are plots of the slope of the linear fits as a function $\omega$ with different system size scaling. Panel (c) clearly shows that the slopes of the different system sizes $L$ but scaled with $L^{1}$ mostly overlap. This implies that the $f$-function flattens as $L^{1}$. As mathematical expressions,
$\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;\xrightarrow{L\to\infty}\;0,\qquad L\;\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;=\;\text{a constant.}$ (3.16)
@@
This monotonically increasing behaviour of $f$-function as a function of $|T|$ actually further reinforces the Kelvin statement, meaning the contribution from this variation is such that it gives extra support in making the change in energy $\Delta E$ to have the same sign as the temperature $T$ or the inverse temperature $\beta$.
@@
$\lambda(t)=2\lambda_{0}\Lambda\,\text{sinc}(\Lambda t)$ (3.17)
[…]
Considering a large enough system size $L$, we can make $\Lambda$ as small as we want without taking it smaller that the random matrix theory (RMT) scale $\sim 1/L^{2}$. We can also ensure that $\Lambda<|T|=1/|\beta|$. Then we can use the expansion of the exponential entropic factor in (3.13). For small $\Lambda$, we consider only the leading order term which gives us
$\Delta E_{n}=\int_{0}^{\Lambda}d\omega\,|\lambda(\omega)|^{2}\,\omega\left(f(\bar{E}+\omega/2,\omega)^{2}-f(\bar{E}-\omega/2,\omega)^{2}\right)\$(3.19)
@@
We want to emphasize that we are working in the small $\omega$ region so this argument is not valid for large $\omega$ region. More crucially, there is no reason why the above expression should be either positive or negative, unlike in subsection 3.1 where we invoked Berry’s conjecture. But it is still a highly desirable result otherwise the chaotic system will violate the second law of thermodynamics in this low energy physics. Note that this leading $\omega^{1}$ term is absent in the thermodynamic limit because the $f$-function goes to the flat-limit of this monotonic behaviour.
Another motivation for the monotonically increasing behaviour of $f$-function as a function $|T|$ comes from the large $\omega$ behaviour of $f(\beta,\omega)$. As a function of $\omega$, it falls exponentially faster than $e^{-|\beta\omega|/4}$. This is indeed a monotonically increasing function of $|T|=1/|\beta|$. [^4] We thank Anatoly Dymarsky for pointing out this result interpreted in terms of the monotonic behaviour of $f(\beta)$. Again we want to emphasize that this exponential fall-off takes into account only $\omega$-dependence in $f(\beta,\omega)$. There could be other $\beta$-dependence even in the large $\omega$ region. But if we take large enough $\omega$, the other $\beta$-dependence cannot trump the exponential monotonic behaviour. So, at least we have this proof of the monotonic behaviour for large enough $\omega$.
@@
This result along with (3.16) constitute ETH-monotonicity. We propose that every quantum chaotic many-body system obeying equivalence of ensembles will have this property.
Note that temperature is an intensive quantity so the monotonically increasing behaviour will remain intact even in the thermodynamic limit. It can be verified from Figure 7 which are plots of $1/f\partial f/\partial\beta$ as a function of $\beta$ for fixed values of $\omega$. Unlike $f(\bar{E})$, $f(\beta)$ does not flatten as the system size increases. But it is not physically relevant because what appeared in (3.13) is $f$-function as a function of $\bar{E}$, not as a function of $T$ or $\beta$.